OpenStudy (anonymous):

Calculate the energy for one mole of photons (an einstein) for light absorbed at 6.80 × 102 nm.

OpenStudy (jfraser):

find the energy of a single photon of that wavelength, then multiply by avogadro's constant\[\nu = \frac{c}{\lambda}\]\[E_{per photon} = h*\nu = h*\frac{c}{\lambda}\]\[E_{mole} = E_{per photon}*\frac{6.022*10^{23} photons}{1 mole \space photons}\]